how to do you solve :
xsquared-4x-2y=0

Found 2 solutions by unlockmath, solver91311:
Answer by unlockmath(1688)

Hello,
This x^2-4x-2y=0is solved by:
Adding 2y to both sides:
x^2-4x=2y
Now divide by 2:
(x^2)/2-2x=y
factor this:
1/2(x)(x-4)
x=0
x=4
RJ
www.math-unlock.com

Answer by solver91311(24713)



You don't. That is a whole set of quadratic equations. The set has infinite elements; one for each possible real value of . Without another relationship defined, you cannot determine which element of the set you are dealing with.

You can apply the quadratic formula to get an expression for the solution:



Where

, , and

You will probably want to specify an interval for such that the roots of the equation will have real values.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism